Short-sales are typically done on margin. Let's assume that you short a stock and your equity is 1/4 of the value of the stock that you shorted. Subsequently, the shorted stock has a return of -8%. What is the return to your equity? Write your answer out to two decimals
Answer #1

Shorted stock initial value=400
Equity value=1/4*400=100
Margin=300
Shorted stock purchase price=400*(1-8%)

Return on equity=400*8%/100=32.00%